If you have an x-intercept = 5 and a y-intercept = -3
1)find the slope and
2) the y-intercept.
How do I do this? I tried setting it up in the equation y2-y1/x2-x1 but I can't seem to get it to work

Found 2 solutions by TimothyLamb, richwmiller:
Answer by TimothyLamb(4379)

x-intercept and y-intercept are two points on the line:
---
given:
x-intercept = (5, 0)
y-intercept = (0, -3)
---
use the two-point form of the linear equation:
---
plug this (the two points given):
5, 0, 0, -3
into the Two-Point form: x1 y1 x2 y2 input box here:
https://sooeet.com/math/linear-equation-solver.php
---
Answer:
slope = 0.6
y-intercept = -3 (already given in the problem statement)
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php

Answer by richwmiller(17219)

we are given the x and y intercept
y intercept means x=0 y=-3 (0,-3) that answers question two
x intercept means y=0 x=5 (5,0)
so we have two points
(5,0) (0,-3)
m=0--3/5-0/=3/5
slope is 3/5 or .6
y-0=3/5*(x-(-3))
y=3/5*x+3